A normal distribution has meu =24 and sigma =3. What is the minimum score needed to be in the top 14% of the distribution>
Answer by ewatrrr(24785)

Population: mu =24 and sigma =3
z = invNorm(.86) = 1.0803 = (X-24)/3
1.0803(3) + 24 = X = 27.2424 (27 not going to quite get You there)
28 minimum score necessary